OpenStudy (anonymous):

A spinner is divided into 10 equal sections numbered 1 through 10. If the arrow is spun twice, what is the probability the first number will be a 2 and the second number will be a 4? A) 0.05 B) 0.125 C) 0.01 D) 0.2

OpenStudy (anonymous):

The probability of two independent events is P(A) * P(B) for event A and B. The probability of hitting either is 1/10, thus the probability of both is 1/100

OpenStudy (cwrw238):

for one spin probability of a 2 is 1/10 as there are 10 equal sections the probability of a 4 is ask 1/10 as these events are independent the probabilities are multiplied so the probability of 2 then 4 = 1/10 * 1/10 = 1/100 or 0.01